OpenStudy (anonymous):

what is y = -2(4x3)8

OpenStudy (bahrom7893):

y = - 2 * 12 * 8 = - 24 * 8 = - 192

OpenStudy (bahrom7893):

please click on Become a fan if I helped, thanks!